# Quantifying Emergent Behaviors in Agent-Based Models using Mean Information Gain

### v1

@@ Abstract
A central challenge in studying complex systems is characterizing and quantifying their emergent behaviors [14]. These behaviors often manifest themselves as distinct visual patterns [13] in one-, two-, or three-dimensional spaces. Efforts to classify these behaviors date back decades [15], and it is customary to use a four-class scheme, distinguishing between the convergent, periodic, chaotic, and complex regimes proposed by Stephen Wolfram [16]. However, much of the field has historically relied on system-specific metrics [17], with limited progress toward generalizable quantitative measures that can capture the essence of emergence across different models. This is where the challenge lies in studying emergent phenomena: we can see the patterns emerging from the interactions within the system, but we still lack a formal, comprehensive approach to fully understand and mathematically describe how these patterns emerge, leaving a critical problem in the study of complex systems unresolved.
In this Letter, we propose a method to quantify differences in system behaviors using a modified conditional entropy metric, the Mean Information Gain (MIG). We show how this metric connects emergence to classical notions of order and disorder in physics. This is illustrated through the application of the MIG to a multi-agent biased random walk in a two-dimensional discrete space, a toy-model capable of reproducing all four behavioral regimes described by Wolfram by just varying the values of its parameters.
@@
Quantifying Emergence.– As a complexity measure, we introduce $\bar{G}$, known as Mean Information Gain or information gain complexity. According to [18], it quantifies the lack of information on other elements of a structure, given that certain properties of the structure are known. It is expressed as
@@
Concluding remarks.– We implemented a multi-agent, rule-based biased random walk model capable of reproducing the four behavioral classes described by S. Wolfram: convergent, periodic, complex, and chaotic. The complex regime displays an emergent behavior in which agents self-organize into a macro-agent that collectively moves resembling a two-dimensional random walk. We employ the Mean Information Gain (MIG), a conditional entropy–based metric that quantifies the lack of information about other elements of a structure, effectively measuring the reduction of the average uncertainty between consecutive spatial configurations. This measure captures the degree of spatial disorder across regimes and provides a natural way to order them from low to high disorder as convergent, periodic, complex, and chaotic. The results confirm that the measure effectively distinguishes ordered behavior from disordered behavior. This provides a reliable quantitative basis for identifying and comparing emergent behaviors that can be readily extended to larger systems. Low MIG in the convergent and periodic regimes reflect their ordered nature, characterized by minimal spatial redistribution of agents. The slight difference observed in terms of variance in the periodic regime probably arises from the presence of multiple clusters oscillating in different directions rather than complete convergence to a single point. In contrast, complex and chaotic regimes produce substantially higher MIG, revealing the metrics’ ability to discern different forms of disorder. The complex regime exhibits coordinated exploration of space, whereas the chaotic regime remains uncoordinated and spatially confined. The numerical similarity observed across both pairs of regimes can be attributed to the spatial confinement imposed by the model and the limited number of occupancy states (occupied or unoccupied), which constrain the resolution of spatial differentiation captured by the MIG.
@@
Future work.– Our toy model exhibits rich and non-trivial behavior, particularly in the complex regime, suggesting that further investigation could provide deeper insights into its dynamics. The emergence of a macro-agent exhibiting a coordinated random walk deserves closer examination. Future work could aim to characterize the type of random walk performed, explore its properties in higher dimensions or alternative spatial topologies, and draw connections to known collective behaviors, such as flocking dynamics. It would also be valuable to identify real-world systems or environments in two-dimensional space that exhibit similar emergent motion. In addition, the model can be extended by treating the vision parameter as continuous rather than binary, allowing smooth transitions between behavioral regimes. Another possible extension involves setting a maximum number of agents per patch to study stacking effects. To better characterize these stacked configurations, the MIG gain could be computed over more complex states, such as the number of agents per cell, providing a finer-grained measurement of spatial complexity and local order.

### v2

@@ Abstract
A central challenge in studying complex systems is characterizing and quantifying their emergent behaviors [14]. These behaviors often manifest themselves as distinct visual patterns [13] in one-, two-, or three-dimensional spaces. Efforts to classify these behaviors date back decades [15], and it is customary to use a four-class scheme, distinguishing between the convergent, periodic, chaotic, and complex regimes proposed by Stephen Wolfram [16]. However, much of the field has historically relied on system-specific metrics [17]. Although previous work has examined both the quantitative [18] and qualitative [11, 19, 20, 21] emergent behavior on cellular automata, it has mainly aimed to classify their rule spaces and computational characteristics with limited progress toward generalizable quantitative measures that can capture the essence of emergence across different models. This is where the challenge lies in studying emergent phenomena: we can see the patterns emerging from the interactions within the system, but we still lack a formal, comprehensive approach to fully understand and mathematically describe how these patterns emerge, leaving a critical problem in the study of complex systems unresolved.
In this Letter, we propose a method to quantify differences in system behaviors using a modified conditional entropy metric, the Mean Information Gain (MIG) derived from the Information Gain defined by [22, 23, 24]. Motivated by Parrott’s use of MIG to quantify ecological complexity [25], we show how this metric connects the emergence to classical notions of order and disorder in physics. This is illustrated through the application of the MIG to a multi-agent biased random walk in a two-dimensional discrete space, a toy-model capable of reproducing all four behavioral regimes described by Wolfram by just varying the values of its parameters.
@@
Quantifying Emergence.– As a complexity measure, we introduce $\bar{G}$ as the MIG. According to [22], it quantifies the lack of information on other elements of a structure, given that certain properties of the structure are known. It is expressed as
@@
Concluding remarks.– We implemented a multi-agent, rule-based biased random walk model capable of reproducing the four behavioral classes described by S. Wolfram: convergent, periodic, complex, and chaotic. The complex regime displays an emergent behavior in which agents self-organize into a macro-agent that collectively moves resembling a two-dimensional random walk. We employ the Mean Information Gain (MIG), a conditional entropy–based metric that quantifies the lack of information about other elements of a structure, effectively measuring the reduction of the average uncertainty between consecutive spatial configurations. This measure captures the degree of spatial disorder across regimes and provides a natural way to order them from low to high disorder as convergent, periodic, complex, and chaotic. The results confirm that the measure effectively distinguishes ordered behavior from disordered behavior. This provides a reliable quantitative basis for identifying and comparing emergent behaviors that can be readily extended to larger systems. Low MIG in the convergent and periodic regimes reflect their ordered nature, characterized by minimal spatial redistribution of agents. The slight difference observed in terms of variance in the periodic regime probably arises from the presence of multiple clusters oscillating in different directions rather than complete convergence to a single point. In contrast, complex and chaotic regimes produce substantially higher MIG, revealing the metrics’ ability to discern different forms of disorder. The complex regime exhibits coordinated exploration of space, whereas the chaotic regime remains uncoordinated and spatially confined. The numerical similarity observed across both pairs of regimes can be attributed to the spatial confinement imposed by the model and the limited number of occupancy states (occupied or unoccupied), which constrain the resolution of spatial differentiation captured by the MIG. Together, these results reinforce the generality of MIG as a measure of structural complexity, supporting its relevance not only in physical systems but also in broader contexts such as ecological organization [25].
@@
Future work.– Our toy model exhibits rich and non-trivial behavior, particularly in the complex regime, suggesting that further investigation could provide deeper insights into its dynamics. The emergence of a macro-agent exhibiting a coordinated random walk deserves closer examination. Future work could aim to characterize the type of random walk performed, explore its properties in higher dimensions or alternative spatial topologies, and draw connections to known collective behaviors, such as flocking dynamics. It would also be valuable to identify real-world systems or environments in two-dimensional space that exhibit similar emergent motion. In addition, the model can be extended by treating the vision parameter as continuous rather than binary, allowing smooth transitions between behavioral regimes. Another possible extension involves setting a maximum number of agents per patch to study stacking effects. To better characterize these stacked configurations, the MIG could be computed over more complex states, such as the number of agents per cell, providing a finer-grained measurement of spatial complexity and local order. Another promising direction is to compare MIG with measures derived from the Kolmogorov algorithmic complexity [26], to test whether it can similarly distinguish the four Wolfram behavioral classes.